in an air motor cylinder the compressed air has an internal energy of 450 kj/kg at the beginning of the expansion and an internal energy of 220kj/kg after expansion.of the work done by air during the expansion under a pressure of 50Pa by 6×10^7cm³ per 10kg.calculate the heat flow to and fro from the cylinder


1
Expert's answer
2021-12-08T10:01:14-0500

"Q=A+\Delta U=\frac{pV}m+U_2-U_1=-200~\frac{kJ}{kg}."